OpenStudy (anonymous):

use the method of substitution to solve the system of linear equations. 3a=2-b 3a=4-b choices: there is one solution. the solution is___. there are infinitely many solutions. there are no solutions. I really need someone's help I only have about 3 minutes more for the computer.

OpenStudy (anonymous):

There is no solution. If 3a=2-b and 3a=4-b then you could say 2-b=4-b If b=2 0=2 which is invalid, so no solution.

OpenStudy (anonymous):

thank you very much, i really appreciate it